OpenStudy (anonymous):

The logistic growth function: f(t)=91965/(1+1513e^-t) What is the limiting size of the population that becomes ill?

OpenStudy (anonymous):

you are being asked for the limit as \(t\to \infty\)

OpenStudy (anonymous):

Knowing that at the beginning, 18 people became ill, and after 2 weeks about 132 people became ill

OpenStudy (anonymous):

Wow you really have my back tonight thank you hahaha

OpenStudy (anonymous):

it is just the number in the numerator

OpenStudy (anonymous):

as \(t\to \infty\) you have \(e^{-t}\to 0\)

OpenStudy (anonymous):

so the denominator gets close to 1 and the whole thing gets close to 91965

OpenStudy (anonymous):

Okay perfect. Thank you very much

OpenStudy (anonymous):

this stuff Knowing that at the beginning, 18 people became ill, and after 2 weeks about 132 people became ill is just extra verbiage to set up the problem yw

OpenStudy (anonymous):

Am I just suppose to assume that e^-t=0?

OpenStudy (anonymous):

in the limit as t gets larger \(e^{-t}=\frac{1}{e^t}\) goes to zero very very fast

OpenStudy (anonymous):

think of it this way \[e^t\] grows very quickly that means \[\frac{1}{e^t}\] has a very big denominator for large values of \(t\) big denominator means the number is very close to zero

OpenStudy (anonymous):

ps don't forget \(e^{-t}\) really means \(\frac{1}{e^t}\)

OpenStudy (anonymous):

right! awesome thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

while youre already here...

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

how would I write this equation in irs equivilent exponential form logv14185=y

OpenStudy (anonymous):

that v is my way of saying like an underscore

OpenStudy (anonymous):

let me guess is it \[\log_{14}(185)=y\]?

OpenStudy (anonymous):

so log(lower exponent 14)185

OpenStudy (anonymous):

yes hahah

OpenStudy (anonymous):

use underscore it works in any case what you need to know is this \[b^y=x\iff \log_b(x)=y\] and you should be able to switch quickly end effortlessly for example \[2^3=8\iff \log_2(8)=3\] and \[10^{-2}=.01\iff \log_{10}(.01)=-2\]

OpenStudy (anonymous):

ooooo I forgot about that property

OpenStudy (anonymous):

so \[\log_{14}(185)=y\iff 14^y=185\]

OpenStudy (anonymous):

they say the same thing, and that is how you are going to solve logarithmic and exponential equations that, plus the change of base formula

OpenStudy (anonymous):

Okay so what would the b, x, and y be for ^5sqrt776=6

OpenStudy (anonymous):

do you need to simplify it first before you determine that?

OpenStudy (anonymous):

Or can you tell just by looking immediately

OpenStudy (anonymous):

i can't really figure out what you wrote give me a second

OpenStudy (anonymous):

I might have written it wrong

OpenStudy (anonymous):

\[\sqrt[5]{776}=6\]

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

i am not sure what you are supposed to do with it it is an equality between numbers you can write it as an exponent though lets see if that is what is needed

OpenStudy (anonymous):

\[\large \sqrt[5]{776}=776^{\frac{1}{5}}\]

OpenStudy (anonymous):

I am asked to give the base, x, and y of that equation

OpenStudy (anonymous):

since \[\large 776^{\frac{1}{5}} =6\] we can also say \[\log_{776}(6)=\frac{1}{5}\]

OpenStudy (anonymous):

so one thing you could say is that the base is \(b=776\) not sure what the \(x\) or \(y\) are supposed to be it is \[\log_b(x)=y\]? in that case \[b=776, x=6,y=\frac{1}{5}\]

OpenStudy (anonymous):

thank you so much you are seriously saving my life tonight lol

OpenStudy (anonymous):

lol not sure if i saved your life, but glad to help

OpenStudy (anonymous):

So if I evaluate log_7(1/^3sqrt7) how would I do that

OpenStudy (anonymous):

so far I have 7^x= 1/(^3sqrt7)

OpenStudy (anonymous):

by sqrt I just mean the symbol. so the cube root of 7 i guess

OpenStudy (anonymous):

or I guess the other way to write that would be (1/(7^1/3))

OpenStudy (anonymous):

so now I have 7^x=1/(7^(1/3))

OpenStudy (anonymous):

\[7^x=\frac{1}{\sqrt[3]{7}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which you wrote as \[7^x=\frac{1}{7^{\frac{1}{3}}}\] but it would help your cause more if you wrote it as \[7^x=7^{-\frac{1}{3}}\]

OpenStudy (anonymous):

You can do that??

OpenStudy (anonymous):

then you would immediately see that \(x=-\frac{1}{3}\)

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

\[b^{-n}=\frac{1}{b^n}\]

OpenStudy (anonymous):

Ah I guess that makes sense

OpenStudy (anonymous):

so \[\large \frac{1}{7^{\frac{1}{3}}}=7^{-\frac{1}{3}}\]

OpenStudy (anonymous):

it is not really a matter of making sense, it is the way you use exponential notation since logarithms are exponents (at least at the introductory level) it is all about knowing exponential notation

OpenStudy (anonymous):

okay i see

OpenStudy (anonymous):

What if I need to evaluate it when it is all in the exponent? Like: 2^(log_2)69

OpenStudy (anonymous):

knowing for example that \[\sqrt[n]{x^m}=x^{\frac{m}{n}}\] and stuff like that

OpenStudy (anonymous):

2^((log_2)69)

OpenStudy (anonymous):

you have \[\large 2^{\log_2(69)}\]?

OpenStudy (anonymous):

I have 2^x=69 but I still have that other 2 to deal with

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok lets go slow because this one is very very easy

OpenStudy (anonymous):

ok sweet haha

OpenStudy (anonymous):

this is not an answer to your question, it is my example to make it clear

OpenStudy (anonymous):

\[2^4=16\] which means \[\log_2(16)=4\] so what is \[2^{\log_2(16)}\]? well it must be \(2^4\) which is \(16\) in other words it is always true that \[\huge 2^{\log_2(\diamondsuit)}=\diamondsuit\]

OpenStudy (anonymous):

the log is an exponent. \(\log_2(69)\) is the the number you raise two to in order to get \(69\)which makes \[\large 2^{\log_2(69)}=69\]

OpenStudy (anonymous):

more generally \[\huge b^{\log_b(x)}=x\]

OpenStudy (anonymous):

so 69?

OpenStudy (anonymous):

for example \[e^{\ln(7)}=7\]

OpenStudy (anonymous):

yes so 69 easy as that

OpenStudy (anonymous):

another way to say this is that the log is the inverse of the exponential

OpenStudy (anonymous):

so \[\log_b(b^x)=x\] and \[b^{\log_b(x)}=x\]

OpenStudy (anonymous):

thank you :) How would I write h(x)=log_5(x)+1

OpenStudy (anonymous):

I am now given multiple equations like these to match up with specific graphs

OpenStudy (anonymous):

i am not sure what you mean "how would i write it"

OpenStudy (anonymous):

\[h(x)=\log_5(x)+1\] looks like a basic log function moved up one unit

OpenStudy (anonymous):

How would I rewrite it I guess.. to graph on a calc. since I cant technically write that in a calculator like that

OpenStudy (anonymous):

oooh i see here take a look at this http://www.wolframalpha.com/input/?i=log_5%28x%29%3D1+domain+.1..20

OpenStudy (anonymous):

you can in fact use a graphing calculator although i would use wolfram i can show you how if you like

OpenStudy (anonymous):

okay can you show me how? And why did you set that equal to 1

OpenStudy (anonymous):

oh damn i put equal one, not plus one my mistake

OpenStudy (anonymous):

I'm thinking for testing purposes I wont have wolfram unfortunately haha

OpenStudy (anonymous):

but that is def helpful for homework

OpenStudy (anonymous):

ok well here is what you can do for a graphing calculator it is always the case that \[\log_b(x)=\frac{\ln(x)}{\ln(b)}\] so for example \[\log_5(x)+1=\frac{\ln(x)}{\ln(5)}+1\]

OpenStudy (anonymous):

that log you do have in your graphing calculator so you can use it if you like but if you have a test it might be hard to match up what you calculator shows you with what is on the test paper so you have to be a bit savvy on how you use it, or else just know something about the graph to begin with

OpenStudy (anonymous):

I'm wolframming these for now and some arent matching up perfectly so I'm just going with what is closest to what it looks like

OpenStudy (anonymous):

you want to post a screen shot or something? i am sure there will be clues to figure it out it is tough with graphs because they all look different you are probably supposed to figure it out some other way

OpenStudy (anonymous):

one clue is where they cross the x or y axis although even that might not be enough

OpenStudy (anonymous):

Yeah okay one sec ill post pics

OpenStudy (anonymous):

A)

OpenStudy (anonymous):

B)

OpenStudy (anonymous):

C)

OpenStudy (anonymous):

before we start is there an equation or are you supposed to match them up

OpenStudy (anonymous):

D)